# jmstate, a Flexible Python Package for Multi-State Joint Modeling

### v1

@@ 8.1 Introduction
The data used in this section originates from the PAQUID cohort [Letenneur et al., 1994], a large prospective population-based study initiated in southwestern France in 1988, aimed at understanding the determinants and trajectories of aging. A subsample of $500$ individuals was followed over a period of up to $20$ years [Proust-Lima et al., 2017], with repeated measures of cognitive and physical health, as well as socio-demographic characteristics. In particular, global cognitive functioning was assessed through the Mini-Mental State Examination (MMSE, see Figure 6), while physical dependency was evaluated using the HIER scale, which classifies subjects into four ordered states (see Figure 7): 0 (no dependency), $1$ (mild dependency), $2$ (moderate dependency), and $3$ (severe dependency). In this work, we focus on the association between cognitive decline and the progression of physical dependency by jointly modeling the longitudinal trajectory of MMSE scores and the transitions between the four HIER states using a joint multi-state model. More specifically, the state of an individual $i$ at a given time $t$ is defined as the highest HIER dependency score between trial entry and $t$. This approach allows us to characterize the dynamic interrelationship between cognitive and physical aging processes within a unified statistical framework.
8.2 Model Specification
Since the inference of the transition specific parameters requires one to have observed those said transitions, the state graph was determined based on the dataset, using the build_buckets function. Moreover, both time and longitudinal values were normalized in [1,1][-1,1]. This technique ensured stable and fast convergence of the optimization process by keeping all parameters on roughly the same scale.
@@ 8.2 Model Specification
Figure 7: Schematic representation of the four ordered HIER states, with possible transitions. HIER quantifies physical dependency (0: no dependency to 33: severe dependency) [Proust-Lima et al., 2017].
We consider an exponential baseline hazard model for transitions, with a clock reset specification. The parameters of these exponential baseline hazards are jointly optimized with the model parameters by passing a list of base_hazard.parameters() to the extra argument of the ModelParams constructor. Moreover, the betas parameter group uses a single shared parameter for all transitions, imposing a particular substructure on the model, where $\forall(k,k^{\prime})\in E,\,\beta^{k^{\prime}\mid k}=\beta$. This is achieved by passing the same tensor for every transition:
# Will be cloned before fitting sharing same memory address shared_beta = torch.zeros(2) init_betas = {key: shared_beta for key in buckets}
In the light of Figure 6, the regression and link functions were taken to be scaled and shifted hyperbolic tangents such that the normalized MMSE score always has a value of $1$ when $t\to$-$\infty$ and is not increasing:
$\begin{cases}h(t,\psi)=g(t,\psi)=\psi_{1}\tanh\left(\frac{\psi_{3}-t}{\psi_{2}}\right)+(1-\psi_{1})\\ \psi=\begin{pmatrix}\sigma(\gamma_{1}+b_{1})\\ \exp(\gamma_{2}+b_{2})\\ \gamma_{3}+b_{3}\end{pmatrix},\quad\sigma(z)=\frac{1}{1+e^{-z}}\end{cases}$
The function $f$ is therefore implemented as:
def f( gamma: torch.Tensor | None, x: torch.Tensor | None, b: torch.Tensor, ): psi = cast(torch.Tensor, gamma) + b psi[..., 0] = psi[..., 0].sigmoid() psi[..., 1] = psi[..., 1].exp() return psi
Please note that type hints are not mandatory, but help with code completion and static type checking; this is the right signature as covariates and populational parameters are passed as optional arguments that may be None.
@@ 8.3.1 Inference
Fitting was performed using $80\%$ of the data, with the remaining $20\%$ used for testing. With no a priori for the covariance matrices, the full option was selected. The model was fitted using the Fit class with a learning rate of $0.5$ and the Adam optimizer. The optimization process was run for until convergence based on the same criterion as the simulation study with a lower tolerance of $5\%$, which was obtained after a little less than $500$ iterations with $10$ parallel chains.
Before displaying the fitted parameters, the Fisher Information Matrix was computed with the ComputeFIM Job factory implementing the method described in Baey et al. [2023], directly taking into account shared parameters:
from jmstate.jobs import ComputeFIM # Returns the metrics for the FIM, and stores them internally model.do( # Uses fitting data if no data is provided job_factories=ComputeFIM(), # No need to wrap in a list max_iterations=100, # Number of samples is n_chains * 100 )
The standard errors were then computed as the square root of the diagonal of the inverse of the Fisher Information Matrix with the model.stderror property. The results are summarized in Table 2.
Parameter Inferred value Standard error Parameter Inferred value Standard error
$\gamma_{1}$ 2.6898 0.0561 $\gamma_{2}$ 0.2626 0.0137
$\gamma_{3}$ 1.3485 0.0025 $\tilde{Q}_{1}$ -0.5428 0.0429
$\tilde{Q}_{2}$ 0.8452 0.1068 $\tilde{Q}_{3}$ 0.5970 0.0448
$\tilde{Q}_{4}$ 0.0041 0.1309 $\tilde{Q}_{5}$ -2.0407 0.1149
$\tilde{Q}_{6}$ 0.2357 0.0458 $\tilde{R}$ 2.1921 0.0135
$\alpha^{3\mid 1}$ -4.1011 1.1296 $\alpha^{1\mid 0}$ 0.5156 2.5595
$\alpha^{2\mid 1}$ -1.5068 1.6333 $\alpha^{2\mid 0}$ -2.9875 2.3893
$\alpha^{3\mid 2}$ -2.7624 0.9352 $\alpha^{3\mid 0}$ -6.2268 4.0880
$\beta_{1}$ -0.0267 0.5393 $\beta_{2}$ 0.1223 0.6959
Table 2: Inferred parameters and their estimated standard deviations (up to 4 decimal places).
@@
The association parameters $\alpha$ being negative, we can infer that cognitive decline is associated with an increase in physical dependency.
@@ 8.3.2 Dynamic Prediction
Using the $n_{\text{test}}=100$ individuals not included in the model fitting, we performed dynamic prediction with the PredictTrajectories Job. For each individual, longitudinal measurements and trajectories were truncated at various times $t$, and predictions were made for future time points $u\geq t$ beyond the truncation. Accuracy was then assessed by comparing the most likely predicted state with the true state at each corresponding future time point $u$, as illustrated in Figure 8.
Formally, the prediction function $\xi_{u}$ is defined by Example 1, and the accuracy measure at time $t$ for predicted states $(s_{i})_{i}$ is defined as:
$\text{accuracy}(u)\coloneqq\frac{1}{n_{\text{test}}}\sum_{i=1}^{n_{\text{test}}}\mathds{1}_{\xi_{u\wedge C_{i}}(\tau_{i},\kappa_{i},\mathcal{T}_{i}^{*})=s_{i}}.$
Although $\mathcal{T}_{i}^{*}$ is unkown, the quantity $\xi_{u\wedge C_{i}}(\tau_{i},\kappa_{i},\mathcal{T}_{i}^{*})$ may be computed based on the censored trajectories. However, incorporating censoring times into prediction functions relies on information that is, in principle, not available at prediction time. Nevertheless, this inclusion is necessary to ensure fairness in accuracy metrics, since individuals who are not censored are generally more likely to have progressed to more advanced states of physical dependency than those who are censored.
@@ 9 Conclusion
We presented jmstate, a Python package for joint modeling of nonlinear longitudinal biomarkers and multi-state survival processes. The framework unifies both components on arbitrary directed graphs, extending classical models to Markov and semi-Markov settings and supporting nonlinear submodels. Inference is implemented with scalable stochastic gradient methods.
@@ Acknowledgments
[…]
@@
$=\lambda_{i}^{S_{i(\ell_{i}+1)}\mid S_{i\ell_{i}}}\big(t\mid\theta,X_{i},T_{i\ell_{i}},b_{i})\exp\Big(-\sum_{s:\,(S_{i\ell_{i}},s)\in E}\Lambda_{i}^{s\mid S_{i\ell_{i}}}\big(t\mid\theta,X_{i},T_{i\ell_{i}},b_{i})\Big).$
Combining both, we retrieve the same joint density as in formula 3.3.
If $t_{i}^{\text{surv}}>$-$\infty$, using the Chasles relation LABEL:rel:chasles, one can also check that both densities are equal. Note that the condtioning only affects the first transition, as the next time $\forall\ell_{i}>0,\,T_{i\ell_{i}}\geq t_{i}^{\text{surv}}\implies\forall\ell_{i}>0,\,T_{i\ell_{i}}\vee t_{i}^{\text{surv}}=T_{i\ell_{i}}$.
[…]
@@
The proof is short and relies on the fact that, using the Semi-Markovian assumption 6:
$p\left((T_{i(l+1)},S_{i(l+1)})_{0\leq l\leq m_{i}-1}\mid\theta,X_{i},b_{i}\right)$ $=\prod_{l=0}^{m_{i}-1}p\left((T_{i(l+1)},S_{i(l+1)})\mid\theta,X_{i},(T_{il^{\prime}},S_{il^{\prime}})_{l^{\prime}\leq l},b_{i}\right),$
$=\prod_{l=0}^{m_{i}-1}p\left((T_{i(l+1)},S_{i(l+1)})\mid\theta,X_{i},(T_{il},S_{il}),b_{i}\right).$
On the other hand, the probability that no additional event is observed between $T_{im_{i}}$ and $C_{i}$ is $1$ if the last state reached is absorbing, and otherwise:
$\mathbb{P}(T_{i(m_{i}+1)}>C_{i}\mid\theta,X_{i},(T_{im_{i}},S_{im_{i}}),b_{i})=\exp\left(-\sum_{s:\,(S_{im_{i}},s)\in E}\Lambda_{i}^{s\mid S_{im_{i}}}(C_{i}\mid\theta,X_{i},T_{im_{i}},b_{i})\right),$

### v2

@@ 7.1 Introduction
The data used in this section originates from the PAQUID cohort [Letenneur1994], a large prospective population-based study initiated in southwestern France in 1988, aimed at understanding the determinants and trajectories of aging. A subsample of $n=500$ individuals was followed over a period of up to $20$ years [Proust-Lima2017], with repeated measures of cognitive and physical health, as well as socio-demographic characteristics. In particular, global cognitive functioning was assessed through the Mini-Mental State Examination (MMSE, see Figure 6), while physical dependency was evaluated using the HIER scale, which classifies subjects into four ordered states (see Figure 7): 0 (no dependency), $1$ (mild dependency), $2$ (moderate dependency), and $3$ (severe dependency). In this work, we focus on the association between cognitive decline and the progression of physical dependency by jointly modeling the longitudinal trajectory of MMSE scores and the transitions between the four HIER states using a joint multi-state model. More specifically, the state of an individual $i$ at a given time $t$ is defined as the highest HIER dependency score between trial entry and $t$. This approach allows us to characterize the dynamic interrelationship between cognitive and physical aging processes within a unified statistical framework.
Both time and longitudinal values were normalized in [1,1][-1,1]. This technique ensured stable and fast convergence of the optimization process by keeping all parameters on roughly the same scale.
@@
Figure 7: Schematic representation of the four ordered HIER states, with possible transitions. HIER quantifies physical dependency (0: no dependency to 33: severe dependency) [Proust-Lima2017].
We consider an exponential baseline hazard model for transitions, with a clock reset specification. The parameters of these exponential baseline hazards are jointly optimized with the model parameters.
Moreover, we impose a particular substructure on the model, where $\forall(k,k^{\prime})\in E,\,\beta^{k^{\prime}\mid k}=\beta$.
In the light of Figure 6, the regression and link functions were taken to be scaled and shifted hyperbolic tangent and its derivative respectively such that the normalized MMSE score always has a value of $1$ when $t\to$-$\infty$ and is not increasing
$\begin{cases}h(t,\psi)=\psi_{1}\tanh\left(\frac{\psi_{3}-t}{\psi_{2}}\right)+(1-\psi_{1}),\\ g(t,\psi)=\frac{\partial}{\partial t}h,\\ \psi=\begin{pmatrix}\sigma(\gamma_{1}+b_{1})\\ \exp(\gamma_{2}+b_{2})\\ \gamma_{3}+b_{3}\end{pmatrix},\quad\sigma(z)=\frac{1}{1+e^{-z}}.\end{cases}$
@@ 7.2.1 Inference
Fitting was performed using $80\%$ of the data, with the remaining $20\%$ used for testing. The model was fitted using the Adam optimizer. The optimization process was run until convergence based on the same criterion as the simulation study with a tolerance of $5\%$, which was obtained after a little less than $500$ iterations with $10$ parallel chains.
Before displaying the fitted parameters, the Fisher Information Matrix was computed, implementing the method described in [baey2023efficient], directly taking into account shared parameters. The standard errors were then computed as the square root of the diagonal of the inverse of the Fisher Information Matrix. The results are summarized in Table 8.
Parameter Inferred value Std. error Parameter Inferred value Std. error
$\gamma_{1}$ 2.1600 0.0470 $\gamma_{2}$ 0.1590 0.0140
$\gamma_{3}$ 1.0010 0.0030 $\tilde{Q}_{1}$ -0.5700 0.0430
$\tilde{Q}_{2}$ 0.9570 0.0960 $\tilde{Q}_{3}$ 0.3820 0.0430
$\tilde{Q}_{4}$ -0.2050 0.1570 $\tilde{Q}_{5}$ -2.3310 0.1380
$\tilde{Q}_{6}$ 0.5220 0.0470 $\tilde{R}$ 2.0830 0.0260
$\alpha^{3\mid 1}$ 0.7220 0.7150 $\alpha^{1\mid 0}$ -0.5400 0.4030
$\alpha^{2\mid 1}$ -1.0030 0.2000 $\alpha^{2\mid 0}$ -0.1210 0.0380
$\alpha^{3\mid 2}$ -0.0480 0.2520 $\alpha^{3\mid 0}$ -0.3070 0.0740
$\beta_{1}$ -0.2990 0.9130 $\beta_{2}$ 0.0970 1.0300
Figure 8: Inferred parameters and their estimated standard deviations. Bold values mean $p<0.05$.
@@
The association parameters $\alpha$ being negative, we can infer that cognitive decline (i.e. rate of change) is associated with an increase in physical dependency.
@@ 7.2.2 Dynamic Prediction
Using the $n_{\text{test}}=100$ individuals not included in the model fitting, we performed dynamic prediction. For each individual, longitudinal measurements and trajectories were truncated at various times $t$, and predictions were made for future time points $u\geq t$ beyond the truncation. Accuracy was then assessed by comparing the most likely predicted state with the true state at each corresponding future time point $u$, as illustrated in Figure 9.
Formally, the prediction function $\chi_{u}^{*}$ is defined by Example 1, and the accuracy measure at time $t$ for predicted states $(s_{i})_{i}$ is defined as
$\text{accuracy}(u)=\frac{1}{n_{\text{test}}}\sum_{i=1}^{n_{\text{test}}}\mathds{1}_{\chi_{u\wedge C_{i}}(\mathcal{T}_{i})=s_{i}}.$
Although $\mathcal{T}_{i}^{*}$ is unkown, the quantity $\chi_{u\wedge C_{i}}(\mathcal{T}_{i})$ may be computed based on the censored trajectories. However, incorporating censoring times into prediction functions relies on information that is, in principle, not available at prediction time. Nevertheless, this inclusion is necessary to ensure fairness in accuracy metrics, since individuals who are not censored are generally more likely to have progressed to more advanced states of physical dependency than those who are censored.
@@ 8 Conclusion
We presented a general likelihood-based framework for joint modeling of nonlinear longitudinal biomarkers and multi-state survival processes. The model unifies both components on arbitrary directed graphs, extending classical joint models to accommodate Markov and semi-Markov structures, recurrent transitions, and nonlinear mixed-effects submodels. This formulation provides a flexible and rigorous foundation for linking longitudinal biomarker trajectories to complex event histories within a single coherent probabilistic model.
@@ Acknowledgments
Conflict of interest
The authors declare that they have no conflicts of interest.
Data availability
The data analyzed in this study are publicly available via the R package timeROC (dataset Paquid).
Appendix A Multi-State Simulation
The simulation of the transition process $\mathcal{T}_{i}^{*}$ conditionally on $b_{i}$, $X_{i}$, and $(T_{i0},S_{i0})$ can be achieved drawing on the semi-Markov property by considering one transition at a time, according to Algorithm 2. The algorithme is similar to Gillespie’s algorithm [gillespie2007stochastic], and can also include a survival condition $T_{i1}\geq t_{i}^{\text{surv}}$ that uses the Chasles relation.
Algorithm 2 Simulation of trajectory $i$
2:
4:
6:  for each $s$ with $(S_{i(\ell_{i}-1)},s)\in E$ do
7:   Draw $T_{i\ell_{i}}^{s}$ such that
8:      $\log\mathbb{P}(T_{i\ell_{i}}^{s}>t)=\int_{T_{i(\ell_{i}-1)}\vee t_{i}^{\text{surv}}}^{t}\lambda_{i}^{s\mid S_{i(\ell_{i}-1)}}(w\mid b_{i},X_{i},T_{i(\ell_{i}-1)},\theta)\,dw,$
9:      $\forall t\geq T_{i(\ell_{i}-1)}\vee u.$
10:  end for
11:
13:  Append $(T_{i\ell_{i}},S_{i\ell_{i}})$ to $\mathcal{T}_{i}$.
14:  $\ell_{i}\leftarrow\ell_{i}+1$.
15:end while
16:
17:if $T_{i(\ell_{i}-1)}>C_{i}$ then
18:  Remove the last pair: $\mathcal{T}_{i}\leftarrow\mathcal{T}_{i}[:-1]$.
19:end if
20:return $\mathcal{T}_{i}$
The proof of the exactness of this algorithm is given below.
@@ Appendix A Multi-State Simulation
$=\lambda_{i}^{S_{i(\ell_{i}+1)}\mid S_{i\ell_{i}}}\big(t\mid\theta,X_{i},T_{i\ell_{i}},b_{i})\exp\Big(-\sum_{s:\,(S_{i\ell_{i}},s)\in E}\Lambda_{i}^{s\mid S_{i\ell_{i}}}\big(t\mid b_{i},X_{i},T_{i\ell_{i}},\theta)\Big).$
Combining both, we retrieve the same joint density.
If $t_{i}^{\text{surv}}>$-$\infty$, one can also check that both densities are equal. Note that the condtioning only affects the first transition, as the next time $\forall\ell_{i}>0,\,T_{i\ell_{i}}\geq t_{i}^{\text{surv}}\implies\forall\ell_{i}>0,\,T_{i\ell_{i}}\vee t_{i}^{\text{surv}}=T_{i\ell_{i}}$.
Appendix B Expression of the Semi-Markov likelihood
@@
The proof is short and relies on the fact that, using the semi-Markov property 2
$p\left((T_{i(\ell+1)},S_{i(\ell+1)})_{\ell\leq m_{i}-1}\mid b_{i},X_{i},\theta\right)$ $=\prod_{\ell=0}^{m_{i}-1}p\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid b_{i},X_{i},(T_{i\ell^{\prime}},S_{i\ell^{\prime}})_{\ell^{\prime}\leq\ell},\theta\right),$
$=\prod_{\ell=0}^{m_{i}-1}p\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid b_{i},X_{i},(T_{i\ell},S_{i\ell}),\theta\right).$
On the other hand, the probability that no additional event is observed between $T_{im_{i}}$ and $C_{i}$ is $1$ if the last state reached is absorbing, and otherwise
$\mathbb{P}(T_{i(m_{i}+1)}>C_{i}\mid\theta,X_{i},(T_{im_{i}},S_{im_{i}}),b_{i})=\exp\left(-\sum_{s:\,(S_{im_{i}},s)\in E}\Lambda_{i}^{s\mid S_{im_{i}}}(C_{i}\mid b_{i},X_{i},T_{im_{i}},\theta)\right),$
@@
Appendix C Stochastic optimization process
(a) $\gamma$
(b) $Q$
(c) $R$
(d) $\alpha^{1\mid 0}$
(e) $\alpha^{2\mid 0}$
(f) $\alpha^{2\mid 1}$
(g) $\beta^{1\mid 0}$
(h) $\beta^{2\mid 0}$
(i) $\beta^{2\mid 1}$